# Projects/QSM24/SFVisco/ViscoSF.tex

### v4

@@ III Results
Figure 5 gives the shear viscosity time function at the lower temperature of $T^{*}=0.80$. Note the change of scale. The viscosity of the quantum liquid is significantly lower than that of the classical liquid over virtually the whole domain. At very small times the quantum viscosity rises more quickly than the classical viscosity before flattening out. This effect was seen consistently in all the simulations, but the quantitative values in the small time regime appear sensitive to the frequency of the application of the dissipative thermostat (see below). The maximum viscosity of the classical liquid is much greater at this temperature than that at the higher temperature of the preceding figure. The root mean square change in separation was 0.54(16) for the quantum liquid and 0.43(7) for the classical liquid. The statistical errors in the average potential energy and pressure were a little smaller when estimated within a run than when estimated from between runs. I conclude from this that the subsystem is a little glassy.
@@
Figure 5 shows the viscosity for the lowest temperature studied, $T^{*}=0.60$. Note the change of scale again. It can be seen that the classical viscosity is substantially larger than the quantum viscosity everywhere except at very short times. The root mean square change in separation was 0.45(14) for the quantum liquid and 0.23(7) for the classical liquid. The statistical errors in the average potential energy and pressure were smaller when estimated within a run than when estimated from between runs, more noticeably so in the quantum case. I conclude from this that the subsystem is somewhat glassy. If this is the case then it is remarkable that the shear viscosity has a finite value, and that it is so low for the quantum liquid.
The results in this paper are based on the computer algorithm described above. Specifically, that the adiabatic transition for the quantum liquid used the shared non-local force, and that the bosons in each momentum state attempted a transition independently one at a time. An alternative is the so-called rigid body model in which the bosons in a momentum state are tied together so that either all or none make the adiabatic transition each time (Attard 2025). The formula for the shear viscosity is the same in both models. The two models should give the same result on average, but, from the fundamental point of view the present one at a time model is arguably preferable because it allows the adiabatic break-up of an occupied state. For the present case of $T^{*}=0.60$ and $\rho^{*}=0.8872$, the present shared non-local force one at a time model gives $\eta^{*}(6)=16.6(59)$, whereas the rigid body all or nothing model gives $\eta^{*}(6)=16.3(75)$.
The present results applied the dissipative block once every 10 time steps, whereas the rigid body results just mentioned applied it once every 5 times steps. Reducing the frequency to once every 25 time steps, (and using runs nearly twice as long) the rigid body model gives $\eta^{*}(6)=19.0(28)$. This is unchanged within the statistical error (although the statistical error itself appears smaller for less frequent applications of the dissipative transition). On very short time scales, $t^{*}\lesssim 0.2$, reducing the frequency of the dissipative transitions causes the viscosity to rise more quickly before flattening out.
@@
Figure 7: Shear viscosity at $t^{*}=6$ on the Lennard-Jones liquid saturation curve. The circles and solid lines are for the quantum liquid $\eta^{\mathrm{qu}}$, the triangles and dashed lines are for the classical liquid $\eta^{\mathrm{cl}}$, and the asterisks and dotted lines are $(1-f_{0}^{\mathrm{qu}})\eta^{\mathrm{cl}}$. The error bars give the 95% confidence level, and the lines are an eye guide.
@@
Figure 7 shows the shear viscosity at $t^{*}=6$ as a function of temperature for the saturated liquid. Based on the extrapolation of fitted quadratics, the viscosity evaluated at this time is relatively close to the maximum value of the viscosity in almost all cases. It can be seen that at higher temperatures the classical and quantum viscosities converge. At the lowest temperature studied the classical viscosity is about 4.5 times larger than the quantum viscosity. Whereas the classical viscosity increases markedly with decreasing temperature, the quantum viscosity is practically constant. This means that the increase in condensation with decreasing temperature in the quantum liquid, Fig. 3, increases the importance of the shared non-local force, and that this reduces the overall shear viscosity to the extent of overcoming the increasing glassiness of the subsystem.
@@
The figure also includes results for $(1-f_{0}^{\mathrm{qu}})\eta^{\mathrm{cl}}(6)$. This viscosity is the analogue of the two-fluid model of superfluidity (Landau 1941, Tisza 1938), namely it is the fraction of uncondensed bosons (ie. those in singly-occupied momentum states) in the quantum liquid times the viscosity in the classical liquid (ie. the viscosity calculated as if all the bosons were in singly-occupied momentum states). We would expect it to have some validity if the viscosity of condensed bosons is zero, and if the actual viscosity is a linear combination of that of the individual components of a two-component mixture. It can be seen that the two-fluid approximation is surprisingly good at lower temperatures. With increasing temperature the quantum and classical viscosities tend toward each other more quickly than the condensation goes to zero. This is the reason that the two-fluid ansatz underestimates the viscosity at the highest temperatures shown.
@@
Figure 8: Component of momentum of a typical boson on a trajectory at $T^{*}=0.60$ and $\rho^{*}=0.8872$, plotted once every 75 time steps. The circles connected by solid lines are for the quantum liquid (offset by $-1$) and the triangles connected by dashed lines are for the classical liquid (offset by $+1$). The dotted lines are eye guides and the dimensionless spacing between momentum states is 0.26.
Figure 8 shows a component of the momentum of a typical boson over time. It can be seen that there is a qualitative difference between the trajectory in the quantum liquid and in the classical liquid. The quantum equations of motion yield a smoother curve, with smaller fluctuations, and noticeable stretches of constant momentum. On this portion of the quantum trajectory, the occupancy of the momentum state that this boson is in ranges up to 26, and averages 5.20. The conclusion is that the shared non-local force in the quantum liquid damps the accelerations experienced by condensed bosons.
@@ IV Conclusion
I say that Bose-Einstein condensation occurs in the low-lying momentum states (Attard 2025). I also say that it is a simplification to divide the liquid into a mixture of two distinct fluids since there are multiple multiply-occupied states. Also, as few as two bosons in the same momentum state experience the effects permutation symmetry. Despite my disdain for the simplicity of the two-fluid model, I must concede that the results in Fig. 7 provide a measure of support for it, at least at low temperatures.
Elsewhere (Attard 2025 §5.3) I have argued that on the far side of the $\lambda$-transition 4He is dominated by momentum loops, which is to say permutations between bosons in the same momentum state. In this case the quantum statistical partition function factorizes into an ideal momentum part and a classical position configuration integral. This predicts that the occupancies of the momentum states are given by ideal statistics, for which the exact analytic results are known (Attard 2025 §2.2, Pathria 1972 §7.1). The present quantum stochastic molecular dynamics results tend to give a smaller occupancy for the ground momentum state, and also for excited momentum states, than that given by ideal statistics. It is possible that this is a finite size effect. Or else it says that interactions play a role in the occupancy and that the proposed factorization (Attard 2025 §5.3) is not entirely accurate. The arguments for the present shared non-local force model relied upon permutations between bosons in neighboring momentum states, not just between those in the same momentum state. In this sense the consequent dynamics are not fully compatible with the factorization model. This matter remains to be clarified.
Although this paper has focussed on the formulation of a computational algorithm for the viscosity of a quantum liquid, it should not be assumed that the molecular equations of motion that are given are merely a numerical technique. The equations of motion yield real trajectories for particles, and there is reason to suppose that expressed in terms of transition probabilities they reflect the underlying reality of nature. This of course has implications for the physical interpretation of quantum mechanics. To which end I suggest
@@ Appendix A Occupation of energy states?
If we add an extra boson and place it in the first state, $N\Rightarrow N+1$, $N_{1}\Rightarrow N_{1}+1$, then the second state must be unaffected. This implies that the eigenfunction of the second state must be independent of the new particle and its position. Hence the energy eigenfunction must factorize into those of the two states,
@@
We conclude that it is meaningless to speak of the occupation of energy states for interacting particles.
@@ Appendix B Momentum eigenfunctions
In the text, and indeed in all of our work on quantum statistical mechanics from the beginning (Attard 2018, 2021), we take the momentum eigenfunctions to be (normalized, unsymmetrized)

### v5

@@ III Results
Figure 4 gives the shear viscosity time function at the lower temperature of $T^{*}=0.80$. Note the change of scale. The viscosity of the quantum liquid is significantly lower than that of the classical liquid over virtually the whole domain. At very small times the quantum viscosity rises more quickly than the classical viscosity before flattening out. This effect was seen consistently in all the simulations, but the quantitative values in the small time regime appear sensitive to the frequency of the application of the dissipative thermostat. The extrapolated maximum viscosity of the classical liquid, $\eta^{\mathrm{cl}}(12)=49(13)$ (compare $\eta^{\mathrm{cl}}(6)=38.0(68)$), is much greater at this temperature than that at the higher temperature of the preceding figure. The root mean square change in separation was 0.45(8) for the quantum liquid and 0.59(9) for the classical liquid. The statistical errors in the average potential energy and pressure in the classical case were less when estimated from within a run as from between runs, suggesting a relatively glassy subsystem.
@@
Figure 5 shows the viscosity for the lowest temperature studied, $T^{*}=0.60$. Note the change of scale again. It can be seen that the classical viscosity is substantially larger than the quantum viscosity everywhere except at very short times. In the quantum case the viscosity time function is relatively flat at the terminus, which suggests that it is close to its maximum, $\eta^{\mathrm{qu}}(6)=21.9(57)$. In the classical case, there is enough curvature and slope to extrapolate to a maximum $\eta^{\mathrm{cl}}(20)=91(53)$, which is substantially larger than the value at the terminus, $\eta^{\mathrm{cl}}(6)=51(17)$.
@@
The root mean square change in separation over a run was 0.31(6) for the quantum liquid and 0.25(5) for the classical liquid. The statistical errors in the average potential energy and pressure were smaller when estimated within a run than between runs. Again the conclusion is that the subsystem is somewhat glassy. If this is the case then it is remarkable that the shear viscosity has a finite value, and that it is so low for the quantum liquid.
@@
Figure 6: Shear viscosity on the Lennard-Jones liquid saturation curve. The circles are for the quantum liquid $\eta^{\mathrm{qu}}(6)$, the triangles are for the classical liquid $\eta^{\mathrm{cl}}_{\mathrm{max}}$. The error bars give the 95% confidence level, and the lines are an eye guide. Inset. Viscosity of the quantum liquid (circles) on a magnified scale compared to two predictions based on the classical viscosity. The asterisks connected by dotted lines are Eq. (3.2), and the plus symbols connected by dash-dot lines are Eq. (3.3). The statistical error of these is larger than that shown for the quantum liquid.
@@
Figure 6 shows the shear viscosity as a function of temperature for the saturated liquid. Based on the flatness of the curves, the quantum viscosity evaluated at $t=6$ is close to its maximum value. The classical viscosity maximum is based on a quadratic best fit extrapolation. It can be seen that at higher temperatures the classical and quantum viscosities converge. At the lowest temperature studied the classical viscosity is about four times larger than the quantum viscosity. Whereas the classical viscosity increases by a factor of eight over the temperature range, the quantum viscosity only increases by a factor of two. The not quite doubling in condensation in the quantum liquid, Fig. 2, is sufficient to cancel almost completely the classical viscosity increase. Evidently condensation reduces the rate of change of momentum via the factor of $1/N_{\bf a}$, and this directly reduces the shear viscosity.
The inset to the figure includes results for
$\eta_{\mathrm{T}}=(1-f_{0}^{\mathrm{qu}})\eta^{\mathrm{cl}}_{\mathrm{max}}.$ (3.2)
@@
This viscosity is the analogue of the two-fluid model of superfluidity (Landau 1941, Tisza 1938), namely it is the fraction of uncondensed bosons (ie. those in singly-occupied momentum states) in the quantum liquid times the viscosity in the classical liquid (ie. the viscosity calculated as if all the bosons were in singly-occupied momentum states). Essentially it assumes that the viscosity of condensed bosons is zero, and it assumes that the actual viscosity is a linear combination of that of the individual components of a two-component mixture.
It can be seen that the two-fluid approximation is surprisingly good. For $T^{*}=0.60$, $\rho^{*}=0.8872$, the quantum viscosity is $\eta^{\mathrm{qu}}(6)=21.9(57)$, and the linear binary mixture result is $\eta_{\mathrm{T}}=24(14)$. For $T^{*}=1.00$, $\rho^{*}=0.7009$, the quantum viscosity is $\eta^{\mathrm{qu}}(6)=11.8(26)$, and the linear binary mixture result is $\eta_{\mathrm{T}}=6.3(19)$.
Of course in the present equations of motion the rate of changed of momentum of condensed bosons is not zero, but is rather reduced by the occupancy of their respective momentum states. For the lowest temperature studied, $T^{*}=0.60$ $\rho^{*}=0.8872$, the average occupancy of momentum states in the quantum case was $\overline{N}_{\mathrm{occ}}=2.455(4)$. For such a small average occupancy it is perhaps surprising that the reduction of the rate of change of momentum is sufficient to reduce the superfluid viscosity by so much. In fact however a plausible model for the viscosity in the condensed regime is
$\eta_{\mathrm{A}}$ $=$ $\frac{1}{\overline{N}_{\mathrm{occ}}^{2}}\eta^{\mathrm{cl}}_{\mathrm{max}}.$ (3.3)
Here $\overline{N}_{\mathrm{occ}}\equiv N/\overline{M}_{\mathrm{occ}}$ is the average occupancy of occupied states, where $\overline{M}_{\mathrm{occ}}$ is the number of occupied states. This factor gives the reduction in the rate of change of the first momentum moment in the force term. This neglects the diffusive (ie. ideal) contribution. This is squared because the viscosity time function is the pair time correlation of the rate of change of the first momentum moment. In the case $T^{*}=0.60$ $\rho^{*}=0.8872$ this formula gives $\eta_{\mathrm{A}}=15.1(89)$, to be compared with the two-fluid model $\eta_{\mathrm{T}}=23.7(139)$ and the actual simulated value $\eta^{\mathrm{qu}}(6)=21.9(57)$. At $T^{*}=1.00$ $\rho^{*}=0.7009$, with $\overline{N}_{\mathrm{occ}}=1.3712(3)$, the respective values are $\eta_{\mathrm{A}}=6.0(18)$, $\eta_{\mathrm{T}}=6.3(19)$, and $\eta^{\mathrm{qu}}(6)=11.8(26)$. It can be seen in the inset to Fig. 6 that the accuracy of this second model is comparable to that of the two-fluid model. This second model has a quantitative justification that does not insist that the viscosity of condensed bosons is zero. This formula explains how the seemingly small average occupancy of condensed bosons is sufficient to cause the reduction in superfluid viscosity comparable to the measured values.
@@
Figure 7: Component of velocity of a typical boson on a trajectory at $T^{*}=0.60$ and $\rho^{*}=0.8872$, plotted once every 75 time steps. The circles are for the quantum liquid (offset by $+1$) and the triangles are for the classical liquid (offset by $-1$). The dotted lines give the zero momentum state for the component. The dimensionless spacing between momentum states is 0.26.
Figure 7 shows a quantum and classical trajectory for a component of momentum of a typical boson. It can be seen that there is a qualitative difference between the trajectory in the quantum liquid and in the classical liquid. The quantum equations of motion yield a smoother curve, with smaller fluctuations, and noticeable stretches of constant momentum. These are correlated with the boson being in a highly occupied momentum state, which are noticeably the low-lying momentum states. On this portion of the quantum trajectory, the occupancy of the momentum state that this boson is in ranges up to 91, and averages 19.4. The conclusion is that the occupancy factor, $1/N_{\bf a}$, damps the accelerations experienced by condensed bosons. It is not hard to imagine that the more frequent changes in momentum evident in the classical liquid dissipate momentum more efficiently and give rise to the usual viscosity of everyday experience.
@@ IV Conclusion
My definition of condensation is different namely that it occurs in the low-lying momentum states (Attard 2025). I also say that it is a simplification to divide the liquid into a mixture of two distinct fluids since there are multiple multiply-occupied states. Also, as few as two bosons in the same momentum state experience the effects permutation symmetry. Despite the simplicity of the two-fluid model, provided that condensation is interpreted as being into the low-lying momentum states, the results in Fig. 6 provide a measure of support for it.
Elsewhere (Attard 2025 §5.3) I have argued that on the far side of the $\lambda$-transition 4He is dominated by momentum loops, which is to say permutations between bosons in the same momentum state. In this case the quantum statistical partition function factorizes into an ideal momentum part and a classical position configuration integral. This predicts that the occupancies of the momentum states are given by ideal statistics, for which the exact analytic results are known (Attard 2025 §2.2, Pathria 1972 §7.1). The present quantum stochastic molecular dynamics results for the occupancy agree with ideal statistics. This is a good test of the computational algorithm and programming.
The present analysis gives an explanation for the molecular origin of superfluidity. The shear viscosity is much reduced or practically zero for condensed bosons because Newton’s second law of motion does not hold: the rate of change of momentum due to an applied force is significantly smaller in the condensed compared to the non-condensed regime. These quantum molecular dynamics equations are necessary for the equilibrium system to evolve at constant entropy, as demanded by fundamental statistical considerations (Attard 2012a Ch. 7), and as confirmed by the thermodynamic analysis of fountain pressure measurements of superfluid helium (Attard 2025 §4.6). In addition, for the case of a subset of bosons changing momentum state, the applied force effective on each individual boson is the shared non-local force, which is reduced from the individual classical force by averaging over the subset to which the boson belongs.
Although this paper has focussed on the formulation of a computational algorithm for the viscosity of a quantum liquid, it should not be assumed that the molecular equations of motion that are given are merely a numerical technique. The equations of motion yield real trajectories for particles, and there is reason to suppose that the transition probability reflects the underlying reality of nature. This of course has implications for the physical interpretation of quantum mechanics and of the universe. To which end I suggest
@@ Appendix A Occupation of energy states?
If I add an extra boson and place it in the first state, $N\Rightarrow N+1$, $N_{1}\Rightarrow N_{1}+1$, then the second state must be unaffected. This implies that the eigenfunction of the second state must be independent of the new particle and its position. Hence the energy eigenfunction must factorize into those of the two states,
@@
I conclude that it is meaningless to speak of the occupation of energy states for interacting particles.
@@ Appendix B Momentum eigenfunctions
In the text, and indeed in all of my work on quantum statistical mechanics from the beginning (Attard 2018, 2021), I take the momentum eigenfunctions to be (normalized, unsymmetrized)